### v1

# Insert In Style: A Zero-Shot Generative Framework for Harmonious Cross-Domain Object Composition

Abstract
Reference-based object composition methods fail when inserting real-world objects into stylized domains. This under-explored problem is currently split between practical “blenders” that lack generative fidelity and “generators” that require impractical, per-subject online finetuning. In this work, we introduce Insert In Style, the first zero-shot generative framework that is both practical and high-fidelity. Our core contribution is a unified framework with two key innovations: (i) a novel multi-stage training protocol that disentangles representations for identity, style, and composition, and (ii) a specialized masked-attention architecture that surgically enforces this disentanglement during generation. This approach prevents the concept interference common in general-purpose, unified-attention models. Our framework is trained on a new 100100k sample dataset, curated from a novel data pipeline. This pipeline couples large-scale generation with a rigorous, two-stage filtering process to ensure both high-fidelity semantic identity and style coherence. Unlike prior work, our model is truly zero-shot and requires no text prompts. We also introduce a new public benchmark for stylized composition. We demonstrate state-of-the-art performance, significantly outperforming existing methods on both identity and style metrics, a result strongly corroborated by user studies.
1 Introduction
Reference-based object composition, focused on the task of inserting a specific object into a scene [8, 22, 47], is a fundamental challenge in computer vision. Recent methods like DreamFuse [15], AnyDoor [6] and IMPRINT [43] have achieved remarkable realism. However, these models are trained almost exclusively on photorealistic data and fail spectacularly when composing objects into stylized domains like paintings, sketches, or digital art—a vast and common use case.
This cross-domain challenge has recently been met by two distinct families of methods. The first, “training-free blenders”, includes pioneers like TF-ICON [27] and, more recently, AIComposer [23]. AIComposer [23] represents the state-of-the-art for this class, cleverly removing the need for the precise text prompts that TF-ICON [27] requires. These methods are fast and practical, but they are fundamentally blenders, not generators. They excel at harmonizing a pasted object but cannot generate a new object natively within the scene, limiting realism.
The second family, “online generators” is represented by Magic Insert [37]. This method achieves high generative fidelity by first finetuning a custom DreamBooth [36] model for a specific object, then performing style injection [7]. However, this quality comes at a prohibitive practical cost. Magic Insert [37] is not zero-shot and requires a slow, computationally expensive, per-subject online finetuning process for every new object. This approach is impractical for real-world, drag-and-drop applications.
Concurrently, general-purpose controllers for DiT models, like OminiControl [45], have proposed unified-attention mechanisms for handling multiple conditions. However, their effectiveness on complex, competing conditions —such as preserving identity while simultaneously transforming style —remains unproven.
The field is thus left with a clear gap: a method that is generative, zero-shot, and architecturally specialized for this competing-condition task.
In this work, we introduce Insert In Style, the first framework to solve this challenge. Our core methodological contribution is two-fold: a novel training protocol and a specialized attention architecture. First, we propose a three-stage training protocol to explicitly disentangle representations: (a) a reference object encoder to learn robust identity, (b) a spatial style encoder to learn generalizable style, and (c) a final composition stage. Second, to surgically enforce this disentanglement at inference time and prevent feature-bleed, we introduce a novel masked-attention mechanism. This specialized architecture stands in direct contrast to general-purpose, unified-attention models and is key to balancing our competing objectives.
Figure 2: Insert In Style generalizes across in-domain and cross-domain tasks. Top (In-domain): The cross-domain specialist method AIComposer [23] incorrectly harmonizes the object. Our method maintains high fidelity, competitive with the in-domain specialist method DreamFuse [15]. Bottom (Cross-domain): DreamFuse [15] fails with a style mismatch, while AIComposer’s [23] harmonization corrupts object fidelity by incorrectly applying background style attributes. Insert In Style uniquely generates a high-fidelity, style-coherent result.
To power this framework, we introduce a 100100k sample training corpus, created via a novel data pipeline that couples large-scale, multi-method generation with a robust, two-stage filtering process. We objectively calibrate our filtering thresholds on a $1,000$ sample, human-annotated validation set, which ensures our final dataset meets a high standard for both identity preservation and style coherence.
Our method is fully zero-shot at inference time. We demonstrate state-of-the-art performance on multiple cross-domain benchmarks, including our new Insert In Style Bench, the largest and most comprehensive public benchmark we introduce for this task. Our method achieves a superior balance of identity preservation and style harmonization, a finding confirmed by extensive evaluation and user studies. Crucially, our model remains competitive on in-domain, photorealistic benchmarks, proving our framework extends a model’s capabilities (see Fig. 2).
Following are the major contributions of this work:
  1. 1.
    A novel generative framework featuring: (i) a three-stage training protocol that learns disentangled encoders for identity and style, and (ii) a masked-attention architecture that prevents feature-bleed between these competing conditions during composition.
  2. 2.
    The largest-scale dataset for this task (100100k samples), curated by a novel, two-stage pipeline that is rigorously calibrated on human annotations to ensure both semantic identity and style coherence. We will make both the dataset and protocol public.
  3. 3.
    A new, diverse, and largest-scale public benchmark, Insert In Style Bench, for evaluating cross-domain object composition, comprising $788$ samples spanning $51$ diverse background styles and $25$ subject categories.
  4. 4.
    State-of-the-art performance, outperforming baselines in quantitative, qualitative, and human evaluations.
2 Related Work
2.1 Generative Object Composition
In-domain Composition.
In-domain composition focuses on realistically inserting an object into a photorealistic scene. Recent methods have excelled at preserving object identity. AnyDoor [6] and MimicBrush [5] use specialized feature extractors, while IMPRINT [43] learns a dedicated identity-preserving representation. Other works leverage DiT [32] architectures, such as DreamFuse [15] and InsertAnything [42], for in-context editing, while ControlCom [53] adds compositional control. While these methods achieve high fidelity in-domain, they are trained almost exclusively on photorealistic data and thus fail to generalize to stylized domains, creating jarring visual mismatches.
Cross-domain Object Composition.
The challenge of cross-domain composition was first addressed by training-free “blender” methods. Pioneers like TF-ICON [27] and its follow-ups, TALE [33] and PrimeComposer [49], manipulate diffusion latents and attention maps to harmonize a pasted object. The state-of-the-art in this class is AIComposer [23], which removes the reliance on precise text prompts. While these methods are fast and practical, they are fundamentally “blend-then-refine” approaches, not true generative models, limiting their realism.
A second family, “online generators”, achieves higher fidelity. Magic Insert [37] represents the state-of-the-art for this approach. It produces high-quality results by finetuning a custom DreamBooth [36] model per-subject. This quality, however, comes at a prohibitive practical cost: Magic Insert [37] is not zero-shot and requires a slow, expensive, online finetuning process for every new object.
Thus, the field faces a clear trade-off: practicality (via AIComposer) versus generative fidelity (via Magic Insert). A framework that is both generative and zero-shot remains a critical open challenge that our work addresses.
2.2 Controllable Diffusion Transformers (DiTs)
The advent of Diffusion Transformers (DiT) [32] marked a shift from traditional UNets, with models like Stable Diffusion $3$ [9] and FLUX.11-dev [19] establishing state-of-the-art performance. This created a need for parameter-efficient adaptation, solved by methods like LoRA [14]. OminiControl [45] emerged as the state-of-the-art general-purpose controller for DiTs, using a “unified attention” to process all conditions jointly.
2.3 Style Transfer
Style transfer is a well-studied field, with methods evolving from early neural approaches to modern [44, 52, 7, 10, 13, 54, 3], high-fidelity diffusion-based techniques. Recent works like CSGO [51] and OmniStyle [50] highlight the critical role of large-scale, high-quality data. However, these methods and their datasets are designed for style transfer, not object insertion. They lack the aligned foreground-reference and object-mask pairs essential for our task. This data gap has been the primary bottleneck for cross-domain composition. Our work is the first to address this by introducing Insert In Style, a large-scale dataset with the precise {foreground reference, stylized scene, foreground mask} triplets required.
Figure 3: Dataset Pipeline. (a) Generation: We create a large-scale, diverse raw corpus by applying a mix of state-of-the-art stylization methods (FLUX.1-Kontext [20], CSGO [51], and CAST [54]). (b) Filtering: Our raw dataset is then refined by our rigorous two-stage filtering process. The Identity Consistency filter prunes samples with semantic drift in the subject region, while the Style Coherence filter removes aesthetic mismatches between the subject region and its surrounding background, together ensuring a high-fidelity dataset.
3 Dataset Generation
Our framework is powered by a new, large-scale dataset. This section details our data curation methodology, which consists of a large-scale corpus generation and our novel filtering pipeline. The entire process is illustrated in Fig. 3. This data-centric approach is the foundation for our model’s zero-shot generative capabilities.
3.1 Data Generation Pipeline
Base Data.
Each sample in our dataset $\mathcal{D}$ originates from a triplet $\{I_{f},I_{c},I_{m}\}$, which includes:
  • a foreground reference image $I_{f}$, serving as the object to be inserted.
  • a composite image $I_{c}$, representing the complete scene with the foreground object already inserted.
  • a corresponding binary mask $I_{m}$, indicating the region of the object in $I_{c}$.
We build upon the DreamFuse dataset [15], which provides high-quality $\{I_{f},I_{c},I_{m}\}$ triplets. Upon inspection, we observed that a subset of these triplets contains a semantic mismatch between the reference $I_{f}$ and the object in the composite image $I_{c}$. To ensure the quality of our base data, we first proactively filter out these mismatched samples using CLIP-based similarity [12] between the foreground reference $I_{f}$ and the masked object region in $I_{c}$. From this cleaned set, we then select foreground references from relevant classes (e.g., object, handheld, animal, pet, and product). This curation process yields our final base dataset, $D_{b}$, of approximately $40,000$ high-quality triplets.
Generation of Stylized Variants.
To generate a training corpus with maximum fidelity and style diversity, we employed a principled, multi-pronged strategy. Our primary generative pipeline is built on FLUX.1-Kontext [20], which we chose for its state-of-the-art performance in structure-preserving stylization [21], [35]. We trained $20$ custom LoRA [14] modules for it on high-quality external style datasets [4] to create variants for popular and complex archetypes, such as Ghibli, Watercolor, and Chinese Ink.
To broaden the style diversity beyond these $20$ archetypes and ensure our model generalizes, we supplemented this pipeline with state-of-the-art reference-based methods. We employed CSGO [51] and CAST [54], which our empirical analysis showed are highly effective at preserving the subject’s spatial integrity and visual identity. We paired these methods with the Style3030k collection [24], a diverse benchmark covering $1,120$ fine-grained style categories.
This combined generation process yields an initial corpus $D_{b}^{sty}$ of approximately 150150k stylized compositions. Each sample in this final dataset consists of a $\{I_{f},I_{c},I_{m},I_{s}\}$ quadruplet, where $I_{s}$ is the stylized composite image.
3.2 Filtering Methods
This large, raw dataset inevitably contains a spectrum of failure cases, including, subject’s semantic identity drift, and local style incoherence. A rigorous filtering stage is therefore essential to curate a high-quality corpus. We propose a novel, two-stage hybrid filtering pipeline to ensure both semantic identity and style coherence.
Filter 1: Identity Consistency Filtering.
After obtaining the raw stylized dataset $D_{b}^{sty}$, it is important to ensure that the identity of the subject remains consistent between the composite image $I_{c}$ and the stylized composite image $I_{s}$. For this, we utilize the CLIP [12] score, which measures semantic similarity, and DINO [30], which measures structural similarity. First, we define an operation $\mathcal{C}(I,I_{m})$ which, for a given image $I$ and mask $I_{m}$, crops the region of $I$ indicated by $I_{m}$. For every sample of $D_{b}^{sty}$, we calculate $S_{\text{clip}}$ and $S_{\text{dino}}$ as follows:
$S_{\text{clip}}$ $=\text{CLIPSim}(\mathcal{C}(I_{s},I_{m}),\mathcal{C}(I_{c},I_{m}))$ (1)
$S_{\text{dino}}$ $=\text{DINO}(\mathcal{C}(I_{s},I_{m}),\mathcal{C}(I_{c},I_{m}))$ (2)
We use a held-out validation set of size $1,000$ from $D_{b}^{sty}$. We manually classify the pairs $I_{c}$, $I_{s}$ based on whether the object identities match. For these same pairs, we calculate the $S_{\text{clip}}$ and $S_{\text{dino}}$ scores. We perform a grid search over the range of the respective scores to find thresholds $\mathcal{T}_{\text{clip}}$ and $\mathcal{T}_{\text{dino}}$ that maximize precision while keeping the rejection rate below a specific limit. The details of this process are presented in the supplementary material. These thresholds are used to filter $D_{b}^{sty}$; we accept only those samples that satisfy $S_{\text{clip}}>\mathcal{T}_{\text{clip}}$ and $S_{\text{dino}}>\mathcal{T}_{\text{dino}}$.
Figure 4: Qualitative samples from our Insert In Style Dataset. Spanning 100100k samples and $1,140$ unique styles, it is the largest-scale corpus for this task. Each <<Subject, Composite, Stylized Composite>> triplet provides the strong, aligned supervision required to train robust, cross-domain insertion models.
Table 1: Comparison with existing style and composition datasets, highlighting the data gap for cross-domain object composition. Our Insert In Style Dataset is the first large-scale corpus to provide all three components essential for this task: a foreground reference ($I_{f}$), a stylized composite image ($I_{c}^{s}$), and a ground-truth object mask ($I_{m}$). Style-focused datasets [50] lack object masks, while insertion-focused datasets [15] lack stylized scenes. Note that “Ref.” is short for “Reference”.
Dataset Venue Task Type Foreground Ref. Composite Image Foreground Ref. Mask Stylized Composite Image # Styles # Samples Style-30K [24] ECCV $2024$ Stylization $1120$ 3030k Wiki-Art [39] arXiv $2015$ Stylization $27$ 5757k ArtBench [25] arXiv $2022$ Stylization $10$ 6060k OmniConsistency [44] NeurIPS $2025$ Stylization $22$ $2600$ OmniStyle [50] CVPR $2025$ Ref. based Stylization (has Style Ref.) $1000$ 150150k DreamFuse [15] ICCV $2025$ Object Insertion - 8484k Insert In Style - Cross-domain Object Insertion $\mathbf{1140}$ $\mathbf{100}$k
Filter 2: Style Coherence Filtering.
While identity-based filtering ensures semantic consistency, it does not guarantee stylistic coherence between the stylized object and its background. We observe that occasionally, stylization disproportionately affects some regions including the subject region, resulting in a perceptually inconsistent stylization between the subject and the rest of the image $I_{s}$. To ensure visual harmony, we use CSD [41], which measures the similarity of style characteristics between two images. We use CSD to calculate style similarity between the subject region and the remaining areas of the image $I_{s}$. For every sample $\{I_{s},I_{m}\}$ from $D_{b}^{sty}$, we calculate the CSD score:
$S_{\text{csd}}=\text{CSD}(\mathcal{C}(I_{s},I_{m}),\mathcal{R}(\mathcal{C}(I_{s},1-I_{m}))),$ (3)
where $\mathcal{R}$ is an operation that copies patches of size $64\times 64$ from retained regions to masked-out regions.The threshold $\mathcal{T}_{\text{csd}}$ for filtering samples is determined in a similar way as described in the identity filtering stage. Using this threshold, we only accept those samples from $D_{b}^{sty}$ that satisfy $S_{\text{csd}}>\mathcal{T}_{\text{csd}}$.
This two-stage filtering process reduces our 150150k generated samples to the final 100100k high-quality, identity-preserving, and style-coherent training corpus. We provide qualitative samples in Fig. 4 and a detailed comparison to existing datasets in Tab. 1 that establishes our corpus as the first and largest to provide the aligned $\{I_{f},I_{s},I_{m}\}$ triplets essential for this task.
4 Proposed Method
The core challenge in cross-domain object composition is to balance two competing objectives: identity preservation (which requires preserving the object’s features) and style harmonization (which requires transforming them). This presents a unique challenge for general-purpose, unified-attention models [45], which process all conditional signals jointly. This joint processing, while powerful, is not explicitly designed to manage competing signals, creating a potential risk of concept interference or feature bleed, where the strong style signal may corrupt identity features, or vice-versa. To solve this, we introduce a novel framework that is a two-fold contribution:
  1. 1.
    A three-stage training protocol that explicitly disentangles these competing concepts by pre-training specialized, independent encoders for identity and style.
  2. 2.
    A specialized masked-attention architecture that surgically enforces this disentanglement during the final composition stage, preventing concept interference.
4.1 Model Architecture
Base Model.
Our architecture extends the FLUX.11-dev [19] framework, a dual-branch DiT [32]. It operates on latent representations: a VAE [17] encodes images into latents $Z_{0}\in\mathbb{R}^{H\times W\times C}$, and a T55 encoder [34] produces text embeddings $Z_{c}\in\mathbb{R}^{L\times D}$. The model is trained as a rectified flow [26] to predict the velocity vector $v_{\theta}$ of a flow matching a linear interpolation between noise $Z_{1}\sim\mathcal{N}(0,I)$ and the target image latent $Z_{0}$. We define the path as $Z_{t}=t\cdot Z_{0}+(1-t)\cdot Z_{1}$, where $t\in[0,1]$. The target velocity is $v^{*}=Z_{0}-Z_{1}$, and the objective is the $L_{2}$ loss:
$\mathcal{L}_{\text{flow}}=\mathbb{E}_{t,\,Z_{0},\,Z_{1}}\left[\left\|v_{\theta}(Z_{t},c,t)-v^{*}\right\|_{2}^{2}\right]$ (4)
Figure 5: Our multi-stage training protocol on a DiT backbone (a). Stages 11 (b) and $2$ (c) are trained in parallel to independently learn object and style encoding. Stage-33 (d) learns composition by assembling these frozen branches, guided by our Structural Mask Attention (e).
Disentangled Conditioning Architecture.
To handle our competing conditions, we extend the FLUX.11-dev [19] architecture with two additional, parameter-efficient conditioning branches. The full model processes four parallel token sequences:
  1. 1.
    Image Latents ($Z_{t}$): The noisy image tokens to be denoised.
  2. 2.
    Text Embeddings ($Z_{c}$): Standard text prompt conditioning.
  3. 3.
    Identity Branch ($Z_{ref}$): A new branch to encode the foreground reference object.
  4. 4.
    Style Branch ($Z_{style}$): A new branch to encode the background style and spatial context.
We initialize the Identity and Style branches with the same architecture and weights as the base FLUX.11-dev [19] image branch. We then insert LoRA [14] adapters into all QKV projections and MLP layers of these two new branches, as well as the main image branch. Only these LoRA parameters are trained.
In each Transformer block, all four token sequences are jointly processed. The QKV matrices for the new conditional branches are computed using the shared weights $W_{I}^{Q,K,V}$ plus their branch-specific LoRA adapters, $\Delta W$:
$\begin{aligned} &Q_{\text{ref}}=Z_{\text{ref}}^{h}(W_{I}^{Q}+\Delta W^{Q}_{\text{ref}}),\\ &K_{\text{ref}}=Z_{\text{ref}}^{h}(W_{I}^{K}+\Delta W^{K}_{\text{ref}}),\\ &V_{\text{ref}}=Z_{\text{ref}}^{h}(W_{I}^{V}+\Delta W^{V}_{\text{ref}})\end{aligned}$ (5)
$\begin{aligned} &Q_{\text{style}}=Z_{\text{style}}^{h}(W_{I}^{Q}+\Delta W^{Q}_{\text{style}}),\\ &K_{\text{style}}=Z_{\text{style}}^{h}(W_{I}^{K}+\Delta W^{K}_{\text{style}}),\\ &V_{\text{style}}=Z_{\text{style}}^{h}(W_{I}^{V}+\Delta W^{V}_{\text{style}})\end{aligned}$ (6)
These are then concatenated with the image ($Q_{t},K_{t},V_{t}$) and text ($Q_{c},K_{c},V_{c}$) tokens for the shared self-attention operation [45], which we detail in Stage $3$.
4.2 Multi-stage Training Protocol
We train the LoRA adapters of our architecture in three distinct stages to incrementally build and align the required representations. Our protocol uses the same frozen, pre-trained FLUX.11-dev [19] model weights as the foundation for three distinct training stages. We first train two specialist encoders independently and in parallel (Stages $1$, 22), then assemble them to train a final compose model (Stage $3$).
Stage 1: Reference Object Encoder.
We first attach a new Identity Branch ($Z_{ref}$) to the frozen, pre-trained FLUX.11-dev [19] base model and train only its LoRA adapters. The model is trained on the Subjects200200K Collection $2$ dataset [45], which contains paired reference objects and their corresponding composed scenes. We use text prompts that implicitly reference the subject (e.g., “a photo of this item on a table”), forcing the model to ground the subject’s identity in the visual features of $Z_{ref}$. The model is trained to reconstruct the full scene using $\mathcal{L}_{\text{flow}}$ (Eq. 4), conditioned on $Z_{t}$, $Z_{ref}$, and $Z_{c}$.
Stage 2: Spatial Style Encoder.
Independently, we attach a new Style Branch ($Z_{style}$) to the same frozen, pre-trained FLUX.11-dev [19] base model and train only its LoRA adapters. This parallel process ensures the style representations are learned completely independently from the identity representations. The model is trained on a diverse corpus of $70,000$ images, including $40,000$ stylized scenes from OmniStyle [50], $15,000$ from StyleBooth [11], and $15,000$ real-world images. We train the model on a style-aware inpainting task. Given an image latent $Z_{i}$ and a binary mask $M$, we define the style context as the unmasked tokens $Z_{\text{style}}=Z_{i}\odot(1-M)$ and the noisy target tokens as $Z_{t}$ (the masked region $Z_{i}\odot M$, noised). The model is trained to denoise $Z_{t}$ conditioned on $Z_{\text{style}}$ and a text prompt $Z_{c}$.
Stage 3: Composition with Masked Attention.
Finally, we assemble the complete compositional model. We load the frozen, pre-trained FLUX.11-dev [19] base model and attach both the pre-trained and frozen Identity Branch from Stage $1$ and the pre-trained and frozen Style Branch from Stage $2$. We now introduce and train a new set of LoRA adapters, $\Delta W_{\theta}$, on the main branch ($Z_{t}$). This stage is trained on our new Insert In Style Dataset containing 100100k samples (see Sec. 3).
To prevent the concept interference between our two competing conditions, we introduce a structural attention mask $\mathcal{M}$. This mask is applied during the shared self-attention calculation to surgically control information flow, as shown in Fig. 5. We define the concatenated query, key, and value matrices as:
$\begin{aligned} Q=[Q_{c};Q_{t};Q_{\text{style}};Q_{\text{ref}}]\\ K=[K_{c};K_{t};K_{\text{style}};K_{\text{ref}}]\\ V=[V_{c};V_{t};V_{\text{style}};V_{\text{ref}}]\end{aligned}$ (7)
The full attention operation then becomes:
$S=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+\mathcal{M}\right)$ (8)
$[Z_{c}^{h+1};Z_{t}^{h+1};Z_{\text{style}}^{h+1};Z_{\text{ref}}^{h+1}]=SV$ (9)
The mask $\mathcal{M}$ (a matrix of 0s and -$\infty$) is configured to enforce two rules: (i) all branches can attend to the text ($Z_{c}$) and image ($Z_{t}$) tokens, but (ii) the Identity Branch ($Z_{ref}$) and Style Branch ($Z_{style}$) are masked from attending to each other. This novel architecture enforces the disentanglement learned in Stages $1$ and $2$, allowing the model to compose the object harmoniously without the style signal bleeding into and corrupting the identity signal, or vice-versa. We ablate this key design choice in Sec. 5.5.
Figure 6: Qualitative comparison with state-of-the-art in-domain and cross-domain baselines. In-domain methods [15, 6] produce jarring style mismatches, failing to generalize. Cross-domain methods [23, 33, 27] corrupt the subject’s identity and fidelity. In contrast, Insert In Style consistently achieves a superior balance, producing results that are both high-fidelity and aesthetically harmonious.
5 Experiments
5.1 Implementation Details
We utilize the PyTorch framework [31] and train all LoRA modules initialized with rank $16$ using Prodigy optimizer [29] with learning rate of $1.0$ for $1$ epoch. All our experiments are conducted on $4$ NVIDIA A100100 GPUs, using a gradient accumulation factor of $2$, resulting in effective batch size of $8$. Both training and inference are performed at spatial resolution of $768\times 768$ pixels.
5.2 Evaluation Benchmarks
AIComposer Benchmark
We benchmark our method on the AIComposer dataset [23], which aggregates $367$ background-foreground pairs and incorporates the $95$ cross-domain samples from the TF-ICON benchmark [27]. The benchmark provides a rigorous test of generalization, featuring a wide array of background styles (e.g., Sketch, Watercolor, Sci-Fi, Pixel Art) and diverse foreground categories (e.g., Animals, Food, Buildings, Cartoon subjects).
Insert In Style-Bench
To rigorously evaluate generalization, we introduce the novel Insert In Style Bench. To our knowledge, it is the largest evaluation benchmark for this task, comprising $788$ challenging pairs. It is specifically designed to test the insertion of photorealistic objects into complex, stylized scenes. It pairs $25$ diverse foreground concepts (e.g., pets, food, toys), sourced from generative models [46] and the Dreambooth dataset [36], with $51$ highly varied backgrounds. To ensure broad stylistic diversity, the backgrounds are meticulously curated from public sources, including Human-Art [16], the Wikiart dataset [39], Kaggle datasets [38, 2, 48, 18, 28], and Pexels [1].
Table 2: Quantitative comparison on AIComposer benchmark dataset. The best results are in bold, and the second best are underlined.
Method Venue CLIP-I $\uparrow$ CSD $\uparrow$ AES $\uparrow$ Overall Mean $\uparrow$
AnyDoor CVPR $2024$ 0.831 0.382 0.611 0.608
DreamFuse ICCV $2025$ 0.784 0.458 0.632 0.625
TF-ICON ICCV $2023$ 0.714 0.438 0.584 0.579
TALE ACMMM $2024$ 0.686 0.495 0.607 0.596
AIComposer ICCV $2025$ 0.774 0.476 0.644 0.631
Insert In Style - 0.779 0.481 0.655 0.638
Table 3: Quantitative comparison on Insert In Style Bench dataset. The best results are in bold, and the second best are underlined.
Method Venue CLIP-I $\uparrow$ CSD $\uparrow$ AES $\uparrow$ Overall Mean $\uparrow$
AnyDoor CVPR $2024$ 0.863 0.318 0.656 0.612
DreamFuse ICCV $2025$ 0.758 0.449 0.681 0.629
TF-ICON ICCV $2023$ 0.687 0.382 0.661 0.577
TALE ACMMM $2024$ 0.671 0.462 0.695 0.609
AIComposer ICCV $2025$ 0.768 0.430 0.692 0.630
Insert In Style - 0.761 0.466 0.697 0.641
5.3 Comparison with Existing Methods
We compare our method against state-of-the-art in-domain object insertion baselines (DreamFuse [15], AnyDoor [6]) and cross-domain composition methods (TF-ICON [27], TALE [33], AIComposer [23]). We present comprehensive qualitative and quantitative results on the benchmarks detailed in Sec. 5.2.
We evaluate all methods across three key aspects: identity preservation, style consistency, and aesthetic quality. We measure identity preservation using CLIP-I [12] between the reference image and the edited region. Style consistency is quantified via CSD [41] between the edited region and the background. Aesthetic quality is measured using a pre-trained Aesthetic Score (AES) model [40]. Crucially, CSD and AES are calculated only when the edit mask (found via pixel differencing and threshold) exceeds $20\%$ of the image. This prevents a known bias where methods that fail to make an edit are unfairly rewarded. We emphasize that relying on any single metric may not fully capture a method’s effectiveness; thus, we introduce an Overall Mean across metrics to ensure a comprehensive evaluation.
Quantitative results are presented in Tab. 2 and Tab. 3 for the AIComposer benchmark [23] and our Insert In Style Bench, respectively. The results in both tables reveal a clear and consistent trade-off in the state-of-the-art: in-domain models (DreamFuse [15], AnyDoor [6]) achieve high identity (CLIP-I) but fail completely on style (low CSD/AES). Conversely, cross-domain methods (AIComposer [23], TALE [33], TF-ICON [27]) achieve better stylization but at a significant cost to object identity (low CLIP-I). Our method is the only one that successfully resolves this dilemma. We simultaneously achieve high identity preservation, strong style consistency, and superior aesthetic quality, outperforming all baselines on this challenging task.
Qualitative results are presented in Fig. 6. These comparisons visually confirm the quantitative findings: in-domain methods like AnyDoor [6] produce jarring stylistic mismatches. Cross-domain methods like AIComposer [23] corrupt the subject’s identity, causing visual artifacts and misapplying background features. Our method consistently achieves a superior balance, producing results that are both high-fidelity and aesthetically harmonious.
Figure 7: User study. In a randomized and blind comparative study, Insert In Style was strongly preferred for “Content Preservation and Style Harmony”, and “Overall Aesthetic Quality”.
5.4 User Study
To evaluate perceptual quality, we conducted a comparative user study $33$ participants. Participants were shown anonymized, randomized results from all methods and asked to select the best output based on two criteria: (a) Content Preservation and Style Harmony, and (b) Overall Aesthetic Quality. The results, presented in Fig. 7, show a strong user preference for our method, which significantly outperformed all baselines in both categories.
5.5 Ablation Study
We validate our complete methodology in Tab. 4 and Fig. 8. The Naïve E2E variant (Row $1$) exhibits a catastrophic failure, inserting random objects (as seen in Fig. 8) and achieving the lowest Overall Mean. This shows our multi-stage protocol is necessary. Tab. 4 reveals a clear Identity-Style trade-off. The w/o Style pre-train variant (Row $3$) achieves the highest CLIP-I but suffers the worst CSD. Conversely, adding the style pre-train without our mask (Row $4$) improves CSD but hurts CLIP-I. This demonstrates the competing objectives and “concept interference” that we aim to solve. Insert In Style (Row $5$), by adding the masked-attention, solves this trade-off: compared to the w/o Masked Attention variant (Row $4$), Insert In Style simultaneously improves CLIP-I, CSD, and AES, achieving the best Overall Mean score.
Figure 8: Qualitative results of the ablation study on our multi-stage training protocol and masked-attention architecture.
Table 4: Ablation on our multi-stage training protocol and masked-attention architecture.
Training Setting CLIP-I $\uparrow$ CSD $\uparrow$ AES $\uparrow$ Overall Mean $\uparrow$
Naïve E2E 0.655 0.455 0.668 0.593
w/o Subject pre-train (Stage $2$ + $3$) 0.726 0.433 0.678 0.612
w/o Style pre-train (Stage $1$ + $3$) 0.778 0.399 0.676 0.618
Full Protocol w/o Masked Attention (Stage $1$ + $2$ + $3$) 0.758 0.452 0.690 0.633
Insert In Style (Full Protocol + Masked Attention) 0.761 0.466 0.697 0.641
6 Conclusion
We introduced Insert In Style, the first zero-shot generative framework for harmonious cross-domain object composition, solving the state-of-the-art’s trade-off between practical “blenders” and impractical “online generators”. Our novel multi-stage training protocol and masked-attention architecture are explicitly designed to manage competing identity and style signals, preventing the “concept interference” common in general-purpose models. Powered by our new 100100k sample dataset, the largest for this task, our method demonstrated state-of-the-art performance across all metrics and was strongly preferred by humans in a user study. We believe our framework, our human-calibrated data pipeline, and our new $788$ sample public benchmark, the largest for this task, open a new avenue for the under-explored task of cross-domain object insertion.